AB bisects CD at E. If CE=2 1/4, calculate CD.
1
Expert's answer
2010-09-15T06:15:03-0400
If AB bisects CD at E it means that CE=ED=2 1/4 and CD=CE+ED=2 1/4+2 1/4=4 1/2
Answer: 4 1/2